# Hierarchical autoregressive neural networks in three-dimensional statistical system

### v1

@@ Abstract
Autoregressive Neural Networks (ANN) have been recently proposed as a mechanism to improve the efficiency of Monte Carlo algorithms for several spin systems. The idea relies on the fact that the total probability of a configuration can be factorized into conditional probabilities of each spin, which in turn can be approximated by a neural network. Once trained, the ANNs can be used to sample configurations from the approximated probability distribution and to evaluate explicitly this probability for a given configuration. It has also been observed that such conditional probabilities give access to information-theoretic observables such as mutual information or entanglement entropy. So far, these methods have been applied to two-dimensional statistical systems or one-dimensional quantum systems. In this paper, we describe a generalization of the hierarchical algorithm to three spatial dimensions and study its performance on the example of the Ising model. We discuss the efficiency of the training and also describe the scaling with the system’s dimensionality by comparing results for two- and three-dimensional Ising models with the same number of spins. Finally, we provide estimates of thermodynamical observables for the three-dimensional Ising model, such as the entropy and free energy in a range of temperatures across the phase transition.
@@ 1 Introduction
The recent idea [1, 2] of employing artificial neural networks in Monte Carlo simulations has given birth to a new class of algorithms that promise to outperform traditional state-of-the-art algorithms, such as the cluster algorithm [3]. In many cases, this objective still remains a future goal because of the problematic scaling with the system size and difficulties in training the neural networks. Hence, the quest for more optimal neural network architectures, see for example [4, 5, 6]. However, several works have observed that such new algorithms can provide access to observables that are difficult to estimate or completely unavailable in traditional Markov Chain Monte Carlo. Examples of such observables include thermodynamic quantities such as the free energy, entropy [7, 8] or information-theoretic observables such as mutual information [9] and quantum entanglement entropy [10]. So far, most applications have been performed for two-dimensional statistical systems or one-dimensional quantum systems.
The hierarchical algorithm proposed in Ref. [11] uses a set of autoregressive neural networks [7] that enable the recursive fixing of all spins on a square lattice. It was devised for two-dimensional spin systems and was benchmarked in the Ising model on a square lattice. The construction relies on the Hammersley-Clifford theorem [12, 13] and the Markov field property of the Ising model, i.e. as long as the interactions are nearest-neighbor only, for any closed contour the probabilities of any spin inside the contour depend conditionally on the values of spins from the boundary. However, the theorem underlying this idea is valid for systems of any number of dimensions.
The goal of this work is to provide a generalization of the hierarchical algorithm to three dimensions and to benchmark its scaling with the system size and the training efficiency. We shall also compare it with the architectures based on convolutional networks. We demonstrate the algorithm using the example of the Ising model.
We consider $N=L^{3}$ spins that are located on a $L\times L\times L$ periodic lattice. The Hamiltonian evaluated on a given configuration of spins, $\mathbf{s}=\{s^{i}\}_{i=1}^{N}$ gives the energy:
$H(\mathbf{s})=-J\sum_{\langle i,j\rangle}s^{i}\,s^{j}\,,$ (1)
@@
$p(\mathbf{s})=\frac{1}{Z}\exp(-\beta H(\mathbf{s}))\,,$ (2)
where $Z$ is the partition function $Z=\sum_{\mathbf{s}}\exp(-\beta H(\mathbf{s}))$ and the sum is performed over all $2^{N}$ configurations.
@@ 2.1 Autoregressive networks
The basic idea of the Variational Autoregressive Networks (VAN) [7] is to represent the probability distribution of the spin configuration (2) as a product of $N-1$ conditional probabilities:
@@
The analytic form of conditional probabilities $p(s^{i}|\cdot)$ is not known, however, one can use neural networks to approximate them. We denote by $q_{\theta}(s^{i}|\cdot)$ the neural approximation, where the subscript $\theta$ indicates the weights of the neural networks. Then,
@@
is the probability distribution of spin configurations provided by the network. There are two crucial properties of the algorithm. First, the network enables a direct sampling of the probability distribution $q_{\theta}$, which is not possible for the target distribution $p$. Second, network training enables $q_{\theta}$ to be a good approximation of $p$, $q_{\theta}\approx p$.
@@
In the VAN algorithm [7], the generation of spin configurations is performed using a single neural network that takes as input the values of the consecutive spins and gives at the output the probabilities of the spins to have value $+1$. The network is evaluated $N$ times and at $i$-th evaluation the (conditional) probability distribution of the $i$-th spin, $q_{\theta}(s^{i}|s^{1},\dots,s^{i-1})$, is calculated based on the values of the $i-1$ spins fixed in the previous steps. Then, the $i$-th spin is fixed using its probability distribution. In principle, such an algorithm can sample any spin system in arbitrary dimensions with any type of interaction.
@@
The architecture of the network needs to follow the functional dependencies of the conditional probabilities, namely $i$-th network’s output must depend only on the first $i-1$ inputs (hence the name ”autoregressive”). To realize this in practice, we rely either on dense layers where some of the weights are multiplied by 0 to ”remove” unwanted connections between the neurons or on the convolutional networks with masked kernels (see Section 3.2 for more details).
@@
The ability of the VAN algorithm to sample from the distribution $q_{\theta}$, where the probability of the spin configuration is explicitly calculated, allows for so-called reverse training using the (reverse) Kullback–Leibler (KL) divergence,
$D_{KL}(q_{\theta}|p)=\sum_{\mathbf{s}}\,q_{\theta}(\mathbf{s})\log\frac{q_{\theta}(\mathbf{s})}{p(\mathbf{s})}=E[\log q_{\theta}(\mathbf{s})-\log p(\mathbf{s})]_{q_{\theta}(\mathbf{s})}.$ (5)
The KL divergence measures the difference between two probability distributions $q$ and $p$: $D_{KL}(q|p)\geq 0$ and $D_{KL}(q|p)=0\Leftrightarrow q=p$ (note however that in general $D_{KL}(q|p)\neq D_{KL}(p|q)$ so $D_{KL}$ is not a distance in the mathematical sense).
One estimates the average on the r.h.s. using some batch of $n$ samples drawn from distribution $q_{\theta}$:
@@
$E[\ldots]_{q_{\theta}}\to\frac{1}{n}\sum_{i=1}^{n}(\ldots),\textrm{ where }\ \mathbf{s}_{i}\sim q_{\theta}.$ (6)
When $p$ is given by (2), one defines:
@@
$\begin{split}\hat{F}_{q}=\frac{1}{n}\sum_{i=1}^{n}\left(\log q_{\theta}(\mathbf{s}_{i})+\beta H(\mathbf{s}_{i})\right),\quad\mathbf{s}_{i}\sim q_{\theta},\end{split}$ (7)
@@
which is minimized during the network training. Minimizing $\hat{F}_{q}$ is equivalent to minimizing (5), as they differ by the additive constant, which is an estimator of the free energy:
@@ 2.2 Neural importance sampling (NIS)
In practice, there is always some difference between $p$ and $q_{\theta}$ as the networks cannot be perfectly trained. This difference introduces a bias when the mean values of observables are calculated. In this manuscript, we shall use the so-called Neural importance sampling (NIS) [2] method to get unbiased results, that is based on the reweighting procedure  [^1] In the alternative approach, Neural Markov Chain Monte Carlo (NMCMC) [1, 2], one uses the accept-reject step on configurations generated by the network.:
$\left\langle{\mathcal{O}}\right\rangle=\frac{1}{N\hat{Z}}\sum_{i=1}^{N}\hat{w}(\mathbf{s}_{i})\mathcal{O}(\mathbf{s}_{i}),\qquad\mathbf{s}_{i}\sim q_{\theta},$ (9)
@@
$\hat{w}(\mathbf{s}_{i})=\frac{e^{-\beta H(\mathbf{s}_{i})}}{q_{\theta}(\mathbf{s}_{i})}$ (10)
@@
$\hat{Z}=\frac{1}{N}\sum_{i=1}^{N}\hat{w}(\mathbf{s}_{i}),\qquad\mathbf{s}_{i}\sim q_{\theta}.$ (11)
It is easy to see that when $q_{\theta}=p$ one has $\hat{w}(\mathbf{s}_{i})=Z$ for all $\mathbf{s}_{i}$ and (9) reduces to the standard average. When $q_{\theta}\neq p$, the importance weights depend on the configuration, and the reweighting (9) compensates for the deviation between those two distributions.
Two notes are in order here. First, it is crucial for the NIS method that $q_{\theta}(\mathbf{s}_{i})$ is known [^2] Some generative models can generate configurations but are unable to provide their probability, for instance GANs. and $>0$ for all generated configurations $\mathbf{s}_{i}$. Second, the definition (11) provides an unbiased estimator of the partition function $Z$, hence NIS gives access to any thermodynamical observable that may be obtained from $Z$. This is the most significant advantage of NIS compared to standard Monte Carlo algorithms, where getting $Z$ is cumbersome.
In the context of the NIS, a natural measure of difference between $q_{\theta}$ and $p$ is the Effective Sample Size ($ESS$)[14, 15]:
$ESS=\frac{\left\langle{\hat{w}}\right\rangle_{q_{\theta}}^{2}}{\left\langle{\hat{w}^{2}}\right\rangle_{q_{\theta}}}\approx\frac{\left(\sum_{i=1}^{N}\hat{w}(\mathbf{s}_{i})\right)^{2}}{N\sum_{i=1}^{N}\hat{w}^{2}(\mathbf{s}_{i})},$ (12)
[…]
@@ 3.1 Hierarchical autoregressive networks
Figure 1: Example of a hierarchical partitioning in 3D for $L=4$. On the first level, the red boundary $B^{(0)}$ is generated with one neural network $\mathcal{N}_{0}$. At the second level of the hierarchy, one neural network of smaller size $\mathcal{N}_{1}$ is used to consecutively fix a set of boundaries shown in green, what makes a network $B^{(1)}$. The remaining empty spins (marked in blue) corresponding to $I^{k}\equiv B^{(2),k}$, $k=1,\dots,8$ have all the neighbors fixed and therefore can be generated from a local Boltzmann distribution with the heatbath algorithm.
@@
In the VAN algorithm, it is necessary to evaluate the neural network $N$ times to fix all $N$ spins. Since all conditional probabilities (4) are provided by a single network, this network must have $N$ input and $N$ output neurons. In the implementation of VAN based on dense neural networks, the number of weights of the network scales as $\sim N^{2}$. Assuming that the whole matrix multiplications are performed (which is usually true, due to the ease of implementation), the cost of generating one configuration scales as $\sim N^{3}$, which is $\sim L^{9}$ for 3D systems. With a different implementation, where for the given conditional probability only the necessary weights get multiplied, this scaling in principle can be reduced to $\sim L^{6}$ (for VAN) or $\sim L^{4}$ (for HAN). However, the gains are not immediately obvious as they depend on the size of the problem and the capacity of the GPU. Such implementation is in progress, but its study is beyond the scope of this paper.
@@
Following the idea of HAN in 2D, we replace a single neural network from the VAN approach by a set of smaller neural networks. In Fig. 1 we show a schematic representation of the spin system of size $4\times 4\times 4$ and demonstrate its partition into recursive subsets. Periodic boundary conditions are imposed, namely, spins in the outer wall interact with the spins in the opposite wall. In the HAN algorithm, we start with the so-called boundary spins, denoted by red dots in Fig. 1 and using a standard autoregressive network we fix their values - we call this the first step of the networks’ hierarchy. This neural network will form the top of the neural networks hierarchy. Thanks to periodic boundary conditions, these three walls provide the boundary that encloses the interior of the system (in Fig. 1 interior is a $3\times 3\times 3$ cube of spins denoted with green and blue dots).
@@
In the second stage of the hierarchy, the second network is used to fix the 3d cross-like structure, denoted by green dots in Fig. 1. The network takes as an input not only the values of spins denoted as green dots, but also the red ones. As the output, it provides conditional probabilities for the green spins. This dependence on the boundary spins was introduced by adding additional input to dense layers keeping the autoregressive dependence of the generated spins, see Fig. 2 in Ref. [11] for the schematic graph.
@@
For $L=4$ the hierarchy ends with 8 blue spins, which are isolated (in the sense that all their neighbors’ values are fixed). The isolated spin can be fixed using the heatbath algorithm, namely one explicitly calculates the probability of the spin $s^{i}$ to be $+1$ from the expression:
[…]
@@
This hierarchy can be straightforwardly extended to larger systems when $L=2^{k},\ k\geq 2$. For $L=8$ at the first stage of the hierarchy one needs to fix spins on 3 walls, each of size $8\times 8$, 169 spins in total because some spins are shared between walls. In the second stage, the 3D cross-like structure has a size of 7 spins in every dimension and divides the spins into 8 cubes each of size $3\times 3\times 3$. At this stage, we can use the same network architecture as that used in the $4\times 4\times 4$ system. Note that the conditional probabilities of those spins have the same functional dependencies on the surrounding spins (according to the Hammersley-Clifford theorem). Therefore, the spins in all 8 cubes (of size $3\times 3\times 3$) can be fixed simultaneously using only one network. In other words, at this stage of hierarchy we have only one network, that fixes the spins in all 8 cubes. The last stage of hierarchy is the isolated spins, that are fixed using the heatbath formula.

### v2

@@ Abstract
Autoregressive Neural Networks (ANN) have been recently proposed as a mechanism to improve the efficiency of Monte Carlo algorithms for several spin systems. The idea relies on the fact that the total probability of a configuration can be factorized into conditional probabilities of each spin, which in turn can be approximated by a neural network. Once trained, the ANNs can be used to sample configurations from the approximated probability distribution and to explicitly evaluate this probability for a given configuration. It has also been observed that such conditional probabilities give access to information-theoretic observables such as mutual information or entanglement entropy. In this paper, we describe the hierarchical autoregressive network (HAN) algorithm in three spatial dimensions and study its performance using the example of the Ising model. We compare HAN with three other autoregressive architectures and the classical Wolff cluster algorithm. Finally, we provide estimates of thermodynamic observables for the three-dimensional Ising model, such as entropy and free energy, in a range of temperatures across the phase transition.
@@ 1 Introduction
The recent idea [1, 2] of employing artificial neural networks in Monte Carlo simulations has given birth to a new class of algorithms that promise to outperform traditional state-of-the-art algorithms, such as the cluster algorithm [3]. This objective remains a future goal because of the problematic scaling with the system size and difficulties in training the neural networks. Hence, the quest for more optimal neural network architectures, see for example [4, 5, 6]. Besides the objective of improved efficiency, several works have observed that such new algorithms can provide access to observables that may be difficult to estimate in traditional Markov Chain Monte Carlo (MCMC). Examples of such observables include thermodynamic quantities such as the free energy, entropy in lattice field theory [7, 8] or information-theoretic observables such as mutual information [9] and quantum entanglement entropy [10, 11]. So far, most applications have been performed for two-dimensional statistical systems or one-dimensional quantum systems, but higher-dimensional systems were also considered – see e.g. [12].
The hierarchical algorithm proposed in Ref. [13] uses a set of autoregressive neural networks [7] that enable a recursive fixing of all spins on a square lattice. It was devised for two-dimensional spin systems and was benchmarked in the Ising model. The construction relies on the Hammersley-Clifford theorem [14, 15] and the Markov field property of the Ising model, i.e., as long as the interactions are nearest-neighbor only, for any closed contour, the probability of any spin inside the contour depends conditionally only on the values of spins from the boundary. However, the theorem underlying this idea is valid for systems in any number of dimensions; for example, in 3 dimensions (3D), a closed surface is needed.
The goal of this work is to provide a generalization of the hierarchical algorithm to three dimensions and to benchmark its scaling with the system size and the training efficiency. We shall also compare it with architectures based on convolutional networks. Again, we demonstrate the algorithm using the example of the Ising model.
We consider $N=L^{3}$ spins that are located on a $L\times L\times L$ periodic lattice. The energy of a configuration of spins, $\mathbf{s}=\{s^{i}\}_{i=1}^{N}$, is given by:
$E(\mathbf{s})=-J\sum_{\langle i,j\rangle}s^{i}\,s^{j}\,$, (1)
@@
$p(\mathbf{s})=\frac{1}{Z}\exp(-\beta E(\mathbf{s}))\,$, (2)
where $\beta$ is an inverse temperature, $\beta=1/T$; $Z$ is the partition function, $Z=\sum_{\mathbf{s}}\exp(-\beta E(\mathbf{s}))$ and the sum is performed over all $2^{N}$ configurations.
@@ 2.1 Autoregressive networks
The basic idea of Variational Autoregressive Networks (VAN) [7] is to represent the probability of a spin configuration (2) as a product of N1N-1 conditional probabilities:
@@
The analytic form of conditional probabilities $p(s^{i}|\cdot)$ is not known; however, one can use neural networks to approximate them. We denote by $q_{\theta}(s^{i}|\cdot)$ the neural approximation, where the subscript $\theta$ indicates the weights of the neural networks. Then,
@@
is the probability distribution of spin configurations provided by the network. There are two crucial properties of the algorithm. First, the network enables a direct sampling of the probability distribution $q_{\theta}$, which is not possible for the target distribution $p$. Second, network training enables $q_{\theta}$ to be a good approximation of pp. [^1] Note that there are other neural samplers, which possess those important properties, e.g., normalizing flows or diffusion models.
@@
In the VAN algorithm [7], the generation of spin configurations is performed using a single neural network that takes as input the values of the consecutive spins and gives at the output the probabilities of the spins to have value +1+1. The network is evaluated $N$ times and at $i$-th evaluation the (conditional) probability distribution of the $i$-th spin, $q_{\theta}(s^{i}|s^{1},\dots,s^{i-1})$, is calculated based on the values of the i1i-1 spins fixed in the previous steps. Then, the $i$-th spin is fixed using its probability distribution. In principle, such an algorithm can sample any spin system in arbitrary dimensions with any type of interaction.
@@
The architecture of the network needs to follow the functional dependencies of the conditional probabilities, namely $i$-th network’s output must depend only on the first i1i-1 inputs (hence the name ”autoregressive”). To realize this in practice, we rely either on dense layers (which is known as the Masked Autoencoder for Distribution Estimation (MADE) [16]) or on convolutional networks. In both cases, some of the weights are multiplied by zero to ”remove” unwanted functional dependencies (see Section 3.2 for more details).
@@
The ability of the VAN algorithm to sample from the distribution $q_{\theta}$, where the probability of a spin configuration is explicitly calculated, allows for the so-called reverse training using the (reverse) Kullback–Leibler (KL) divergence,
$D_{KL}(q_{\theta}|p)=\sum_{\mathbf{s}}\,q_{\theta}(\mathbf{s})\log\frac{q_{\theta}(\mathbf{s})}{p(\mathbf{s})}=\left\langle{\log q_{\theta}(\mathbf{s})-\log p(\mathbf{s})}\right\rangle_{q_{\theta}},$ (5)
where $\left\langle{\ldots}\right\rangle_{q_{\theta}}:=\sum_{\mathbf{s}}q_{\theta}(\mathbf{s})(\ldots)$ is a statistical average over the distribution $q_{\theta}$. The KL divergence measures the difference between two probability distributions $q$ and $p$: $D_{KL}(q|p)\geq 0$ and $D_{KL}(q|p)=0\Leftrightarrow q=p$ (note, however, that in general $D_{KL}(q|p)\neq D_{KL}(p|q)$ so $D_{KL}$ is not a distance in the mathematical sense). In a typical situation in Machine Learning, where data are easily available, one usually applies forward training, where $D_{KL}(p|q_{\theta})$ is minimized. This is also possible in the context of statistical physics or lattice field theory; however, it requires configurations generated from $p$ using, for example, some MCMC algorithm. One estimates the average on the r.h.s. of (5) using a batch of $n$ samples drawn from distribution $q_{\theta}$:
@@
$\left\langle{\ldots}\right\rangle_{q_{\theta}}\to\frac{1}{n}\sum_{i=1}^{n}(\ldots),\textrm{ where }\ \mathbf{s}_{i}\sim q_{\theta}.$ (6)
If $p$ is given by (2), one defines:
@@
$\begin{split}\beta\hat{F}_{q}=\frac{1}{n}\sum_{i=1}^{n}\left(\log q_{\theta}(\mathbf{s}_{i})+\beta E(\mathbf{s}_{i})\right),\quad\mathbf{s}_{i}\sim q_{\theta},\end{split}$ (7)
@@
which is minimized during the network training. Minimizing $\beta\hat{F}_{q}$ is equivalent to minimizing (5), as $D_{KL}(q_{\theta}|p)=\beta F_{q}-\beta F$, where $F_{q}=\left\langle{\log q_{\theta}+\beta E}\right\rangle_{q_{\theta}}/\beta$ and the additive constant is the free energy:
@@ 2.2 Neural importance sampling (NIS)
In practice, there is always some difference between $p$ and $q_{\theta}$ as the networks cannot be perfectly trained. This difference may introduce a bias when the mean values of the observables are calculated. In this manuscript, we shall use the so-called Neural importance sampling (NIS) [2] method to get unbiased results, which is based on the reweighting procedure. [^2] In the alternative approach, Neural Markov Chain Monte Carlo (NMCMC) [1, 2], one uses the accept-reject step on the configurations generated by the network. For some observable $\mathcal{O}(\mathbf{s})$, the expectation value is given by:
$\left\langle{\mathcal{O}}\right\rangle=\frac{1}{{\cal N}\hat{Z}}\sum_{i=1}^{\cal N}\hat{w}(\mathbf{s}_{i})\mathcal{O}(\mathbf{s}_{i}),\qquad\mathbf{s}_{i}\sim q_{\theta},$ (9)
@@
$\hat{w}(\mathbf{s}_{i})=\frac{e^{-\beta E(\mathbf{s}_{i})}}{q_{\theta}(\mathbf{s}_{i})},$ (10)
@@
$\hat{Z}=\frac{1}{{\cal N}}\sum_{i=1}^{\cal N}\hat{w}(\mathbf{s}_{i}),\qquad\mathbf{s}_{i}\sim q_{\theta}.$ (11)
It is easy to see that when $q_{\theta}=p$ one has $\hat{w}(\mathbf{s}_{i})=Z$ for all $\mathbf{s}_{i}$ and (9) reduces to the standard average. When $q_{\theta}\neq p$, the importance weights depend on the configuration, and the reweighting (9) compensates for the deviation between these two distributions.
Two notes are in order here. First, it is crucial for the NIS method that $q_{\theta}(\mathbf{s}_{i})$ is known [^3] Some generative models can generate configurations but are unable to provide their probability, for instance Generative Adversarial Networks (GANs). and is nonzero for all configurations in support of $p$. Second, the definition (11) provides an unbiased estimator of the partition function $Z$; hence, NIS gives access to any thermodynamic observable that may be obtained from $Z$ – see the end of section 4.2. This is the most significant advantage of NIS compared to standard Monte Carlo algorithms, where getting $Z$ is cumbersome.
In the context of NIS, a useful measure of the difference between $q_{\theta}$ and $p$ is the Effective Sample Size (ESSESS)[17, 18]:
$ESS=\frac{\left\langle{\hat{w}}\right\rangle_{q_{\theta}}^{2}}{\left\langle{\hat{w}^{2}}\right\rangle_{q_{\theta}}}\approx\frac{\left(\sum_{i=1}^{\cal N}\hat{w}(\mathbf{s}_{i})\right)^{2}}{{\cal N}\sum_{i=1}^{\cal N}\hat{w}^{2}(\mathbf{s}_{i})},$ (12)
where $\approx$ denotes the approximation of a finite batch size of samples drawn from $q_{\theta}$. Values of $ESS$ are in the range $(0,1]$ and for $q_{\theta}=p$ one has ESS=1ESS=1.444Note that $ESS=1$ does not mean that $q_{\theta}=p$ since the mode collapse may occur, see discussion at the end of section 4.2. The ESS is a measure of how effective the sampler is, namely the statistical uncertainty of the observable obtained using NIS is $\sigma_{stat}\propto 1/\sqrt{{\cal N}\cdot ESS}$, where $\cal N$ is a number of generated samples.
@@ 3.1 Hierarchical autoregressive networks
Figure 1: Cube on the left part of the Figure: a hierarchical partitioning in 3D for $L=4$. The full system of $4\times 4\times 4$ spins is divided into 3 subsets (denoted with the colors). Central and right part of the Figure: for better visibility of the 3D system, we also show four slices of the $4\times 4\times 4$ cube. Spins generated with the networks at a given step of hierarchy are denoted with colors (see the text for details): red (first step and network), green (second step and network), blue (third step – heat-bath algorithm).
@@
In the VAN algorithm, it is necessary to evaluate the neural network $N$ times to fix all $N$ spins. Since all conditional probabilities (4) are provided by a single network, this network must have $N$ input and $N$ output neurons. In the implementation of VAN based on dense neural networks (MADE), the number of weights of the network scales as $\sim N^{2}$. Assuming that the whole matrix multiplications are performed (which is usually true, due to the ease of implementation), the cost of generating one configuration scales as $\sim N^{3}$, which is $\sim L^{9}$ for 3D systems. Note that this naive counting does not take into account the numerical cost of handling data, etc., which can significantly modify the scaling at small $L$.
@@
The unfavorable scaling of the numerical cost of VAN, coming from the evaluation of a large network multiple times, is an inspiration for the development of an improved algorithm, called Hierarchical Autoregressive Networks (HAN). In HAN, one replaces a single neural network from the VAN approach with a set of smaller neural networks. In Fig. 1 we show a schematic representation of the spin system of size $4\times 4\times 4$ and demonstrate, using colors, its partition into recursive subsets (cube on the left). The right part of Fig. 1 presents 4 slices of the 3D cube (presented for better visibility). Periodic boundary conditions are imposed, namely, spins in the outer wall interact with the spins in the opposite wall.
We start by explaining the HAN algorithm for the case $L=4$ (shown in Fig. 1) and later discuss larger systems. For $L=4$ HAN 3D consists of the following steps:
  1. 1.
    We start generation with the so-called boundary spins, denoted by red dots in Fig. 1. For this, we are using a standard autoregressive network - we call this the first step of the network’s hierarchy. Due to periodic boundary conditions, these three walls provide the boundary that encloses the interior of the system (in Fig. 1 interior is a $3\times 3\times 3$ cube of spins denoted with green and blue dots).
  2. 2.
@@
    In the second stage of the hierarchy, the second network is used to fix the 3d cross-like structure, denoted by green dots in Fig. 1. The network takes as input not only the values of spins denoted as green dots but also the red ones. As the output, it provides conditional probabilities for the green spins. This dependence on the boundary spins was introduced by adding additional input to dense layers, keeping the autoregressive dependence of the generated spins, see Fig. 2 in Ref. [13] for the schematic graph.
  3. 3.
@@
    For $L=4$, the hierarchy ends with 8 blue spins, which are isolated (in the sense that all their neighbors’ values are fixed). The isolated spin can be fixed using the heatbath algorithm, namely, one explicitly calculates the probability of the spin $s^{i}$ to be +1+1 from the expression:
    $q(s^{i}=+1,n(s^{i}))=\left[1+\exp\left(-\,2\beta\sum_{\ j\in n(s^{i})}s^{j}\right)\right]^{-1},$ (13)
@@
For $L=8$ at the first stage of the hierarchy, one needs to fix spins on 3 walls, each of size $8\times 8$, 169 spins in total, because some spins are shared between walls. In the second stage, the 3D cross-like structure has a size of 7 spins in every dimension and divides the spins into 8 cubes, each of size $3\times 3\times 3$. At the third step, we can use the same network architecture as that used in the $4\times 4\times 4$ system (green spins in Fig. 1). Note that the conditional probabilities of those spins have the same functional dependencies on the surrounding spins - this is a consequence of the Hammersley-Clifford theorem (see Section 1). Therefore, the spins in all 8 cubes (of size $3\times 3\times 3$) can be fixed simultaneously using only one network. In other words, at the third step of the hierarchy in $L=8$, we have one network that fixes the cross-like structures (green dots in Fig. 1) in eight cubes in parallel. The last stage of hierarchy is the isolated spins, which are fixed using the heat-bath step with probabilities (13).